# Scaling Test-time Compute for Low-resource Languages: Multilingual Reasoning in LLMs

### v1

@@ 4.1 Experiment settings
Figure 4: Representation retrieval accuraccy between Left: questions, and Right: questions and reasoning traces of the same questions in different languages.
Table 3: Ablation study of fine-tuning on French - a medium resource language, and benchmark across English and Fench reasoning datasets. Bold scores indicate the best performance per benchmark.
@@
r1-distill-Llama-8B Base 43.33 79.6 49.74 54.8
r1-distill-Llama-8B Native CoT Training [1, 13] 45.00 82.0 31.90 61.2
@@
DeepHermes-3-Llama-3-8B Base 1.67 38.8 3.69 21.6
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 3.33 70.8 6.26 44.8
@@
Model Setting AIME2024 (en) MSGM (en) MSGM (zh)
r1-distill-Llama-8B Base 43.33 79.6 65.2
r1-distill-Llama-8B Native CoT Training [1, 13] 46.33 92.4 81.6
DeepHermes-3-Llama-3-8B Base 1.67 85.2 50.8
DeepHermes-3-Llama-3-8B Native CoT Training [1, 13] 11.67 83.2 61.6
4.2 Results and analysis
Reasoning capability can be effectively transferred to solve problems in low-resource language. Table 1 presents the performance of models trained with our approach (English-Pivoted CoT Training), compared to other baselines on English and Irish benchmarks. Our method demonstrates a strong improvement over the baselines. On the Irish version of AIME2024, while all baselines perform poorly, our approach obtains a clear gap of upto 28.33%, in the case of the r1-distill-Llama-8B baseline. More specifically, we achieve accuracies of 35.00% and 10.00% when fine-tuned on r1-distill-Llama-8B and DeepHermes-3-Llama-3-8B, respectively, matching performance on the English version. On LC2024, which comprises relatively less challenging math problems than AIME2024, although the r1-distill-Llama-8B baseline achieves an acceptable accuracy of 63.64%, our approach surpasses it by a 10% margin (73.33% vs. 63.64%). Further inspection shows that both baseline models generate all chain-of-thoughts and final answers in English, regardless of prompting in which languages. This default behavior aligns with the prompting approaches in [21, 23], where they prompt engineering the model to reason in English to obtain higher performances.
Furthermore, the TIES-merging and Layer Swap approaches fail and even unable to correctly answer any of the questions on AIME2024, likely due to significant parameter discrepancies when merging the base model trained on English reasoning tasks with a model trained on general instruction tasks in the low-resource language. These differences likely cause interference between subsets of merged parameters. Another interesting observation is that Native CoT Training, training entirely in the target language for both the chain-of-thought and final answer improved performance on AIME2024 but degraded it on LC2024 (63.64% to 37.14% for r1-distill-Llama-8B), potentially due to reduced reasoning capability when forced to reason in a low-resource language.
Improved language understanding capability. To further investigate whether our approach also enhances language understanding or simply overfits to mathematical phrases in the target language, we split LC2024 into its two original sections: Concepts and Skills and Contexts and Applications. The latter, comprised of more contextual and real-world application descriptions, requires deeper language understanding in Irish. As shown in Table 2, Native CoT training significantly reduces the model’s reasoning capability, as evidenced by the drop from 45.45 to 13.64 on Concepts and Skills for DeepHermes-3-Llama-3-8B, while maintaining the same level of performance for r1-distill-Llama-8B. This supports the idea that forcing the model to reason in the low-resource language affects its ability to generate effective reasoning traces. In contrast, our approach yields gains substantially in Contexts and Applications split — indicating that not only does English-Pivoted CoT Training enable effective reasoning in low-resource languages, but it also improves overall language understanding.
@@ 4.2 Results and analysis
Figure 5: Changes in correction rate from baseline model (r1-distill-Llama-8B) when fine-tuned with our proposed approach across 3 benchmark datasets on 3 languages, from low (Irish) to medium (French) to high (Chinese) resource.
Diverged reasoning traces when forced to reason in different languages. Furthermore, we analyze the model’s representations across languages by computing and comparing the average token representations for both the input alone and the input plus the generated reasoning trace on AIME2024. We measure the retrieval accuracy for matching representations across languages with either questions or questions concatenating with reasoning traces on AIME2024. The retrieval task involves finding the closest English sample given an Irish sample representation, using average token embeddings at any layer. Figure 4 visualizes the accuracies across layers. While models trained solely in the target language (Native CoT Training) show strong alignment for questions (even upto 100% at middle layers), their alignment significantly drops (to highest of around 35% at 2 layers) when a chain-of-thought is included — highlighting discrepancies when the model is forced to reason in a different language than the final output.
This also suggests that question alignment [36, 14, 9], either through parallel sample fine-tuning or prompting, is not enough. Figure 4 shows that while the alignment between questions in different languages is inherently high, this can still lead to really diverged reasonings and answers. Figure 1 depicts an example, where the whole embedding of the questions are similar, a small difference (e.g., mixing between independence and dependence) can lead to a totally different interpretation.
In contrast, our approach - English-Pivoted CoT Training, which trains the model to understand the target language while reasoning in English, achieves the strongest alignment between the reasoning traces (almost close to 100%), demonstrating a more consistent internal representation.
Generalizability to medium- and high-resource languages. We apply our proposed approach to Chinese and French languages, which can be considered medium to high-resource languages [33, 3]. The results, present in Table 3 and Table 4, indicate that our approach are generalizable to other languages. First, the baseline models perform better compared to on the extremely low-resource language, Irish, with a score of 54.8% and 65.2% accuracy on MSGM (French and Chinese versions, respectively) for r1-distill-Llama-8B. Nevertheless, we see a sustainable improvement on French when leveraging our approach, of 83.2% to 54.8% on MGSM (French version). However, on Chinese, as the model already performs well, our approach does not lead to much improvement. Furthermore, qualitative analysis of the reasoning traces of the base models shows that the models actually perform reasoning with CoT in Chinese, compared to English, when prompted with Chinese problems versus problems in low and medium-resource languages. Therefore, forcing the model to reason in English can create an interference. This also explains why further training on target-language reasoning traces (Native CoT Training) can improve the performance (81.6% compared to 70.0% of OURS and 65.2% of baseline).
@@
Figure 5 shows the correction rate of fine-tuned model with our approach, compared to the baseline r1-distill-Llama-8B. More specifically, the percentage of the new model corrects an incorrect answer by the base model and vice versa. We can see that on MSGM (zh), the correct $\rightarrow$ incorrect rate is the highest, highlighting conflicts and diverged solutions by forcing the model to think in English for Chinese problems when they already have a tendency to think in Chinese for Chinese problems. Nevertheless, our method still achieves an average improvement of 14.30% across the 3 resource regimes (low, medium, and high-resources), demonstrating its generalizability.
5 Conclusion
@@
In this work, we propose a novel training approach, named English-Pivoted CoT Training, to effectively transfer the reasoning capabilities of large language models to low-resource languages. Through training to explicitly aligning reasoning processes across languages by forcing CoT traces to be in the dominant language, our approach achieves significant performance gains (up to 28.33%) over existing multilingual reasoning techniques, and is generalizable to other resource regimes (medium and high-resource languages). Furthermore, our analysis provide insights into multilingual LLM behavior, particularly the benefit of separating language understanding from reasoning. By contributing the first-ever Irish mathematical reasoning benchmark (LC2024), we also aim to support future research in multilingual reasoning. Future directions include exploring the generalizability of our approach to other low-resource languages and tasks beyond mathematics, as well as investigating strategies for further improving cross-lingual reasoning capabilities.
{ack}
This research work has emanated from research conducted with financial support from Science Foundation Ireland under Grant 12/RC/2289-P2 and 18/CRT/6223.

### v2

@@ Introduction
Figure 4: Representation retrieval accuracy between Left: questions, and Right: questions and generated CoT traces of the same questions in different languages.
Results and analysis
Reasoning capability can be effectively transferred to low-resource languages. Table 1 presents the performance of models trained with our approach (English-Pivoted CoT Training) compared to baselines on English and Irish benchmarks. Our method demonstrates a strong improvement: on the Irish version of AIME2024, while all baselines perform poorly, our approach obtains a clear gap of upto 28.33% on r1-distill-Llama-8B baseline. More specifically, we achieve accuracies of 35.00% and 10.00% when fine-tuned on r1-distill-Llama-8B and DeepHermes-3-Llama-3-8B, respectively, matching performance on the English version. On LC2024, which is relatively less challenging than AIME2024, although r1-distill-Llama-8B achieves an acceptable accuracy of 63.64%, our approach surpasses it by a 10% margin (73.33%). Further inspection shows that both baseline models generate all CoTs and final answers in English, regardless of prompting in which languages. This default behavior aligns with the English prompting approaches in (qin-etal-2023-cross; shi2023language). Interestingly, performances on LC2024 (authored by Irish speakers) drops compared to the translated Irish AIME2024 for all baseline approaches. We hypothesize the model relies on familiar math notation (same across all languages) to infer the task and then applies English reasoning, whereas interventions such as prompting, CoT fine‑tuning or model merging disrupt this implicit strategy, causing parameter interference in reasoning in an extremely low-resource language.
Improved language understanding capability. To investigate whether our approach also enhances language understanding or simply overfits to mathematical phrases, we split LC2024 into its two original sections: concepts & skills and contexts & applications - with more contextual and real-world application descriptions, requires deeper Irish language understanding. Due to space limitation, we present the results of our method and 2 strongest baselines, with more results in the Appendix. As shown in Table 2, Native CoT training reduces the model’s performance for DeepHermes-3-Llama-3-8B from 45.45 to 13.64 on concepts & skills, while maintaining the same performance on r1-distill-Llama-8B. This shows that training to force the model to reason in low-resource languages affects its ability to generate effective CoTs. In contrast, our approach yields substantially gain in contexts & applications, indicating that not only does it enable effective reasoning in low-resource languages but also improves language understanding.
Training Setting Prompting Setting AIME2024 LC2024 - concepts&skills LC2024 - contexts&applications
Mixed CoT Lang. Irish CoT 33.33 3.33 54.55
Mixed CoT Lang. English CoT 33.33 30.00 61.82
Two-stage Irish CoT 40.00 6.67 45.45
Two-stage English CoT 33.33 33.33 58.18
Table 3: Irish task performance by CoT language for r1-distill-Llama-8B trained on English and Irish CoTs.
Diverged reasoning traces when forced to reason in different languages. We assess cross‑language alignment on AIME2024 by retrieving English embeddings for Irish questions (with and without CoTs) using average token representations across layers. As illustrated in Figure 4, target‑language‑only models show near‑perfect question‑only alignment (up to 100%), but this falls to around 35% once CoTs are included, revealing divergent internal representations. Figure 1 illustrates how minor embedding shifts (e.g., confusing “independence” and “dependence”) can alter interpretation. In contrast, English‑Pivoted CoT Training maintains almost 100% alignment for both questions and reasoning traces, demonstrating stable, language‑agnostic embeddings.
Additionally, we experience with controlling the CoT language at test time. We introduce 2 additional training settings: 1) Mixed CoT Language: 50% of CoTs in Irish, 50% in English, 2) Two‑Stage: first stage uses English‑Pivoted CoT to align comprehension, then a second stage fine‑tunes on native Irish CoTs to refine reasoning in the target language. We prepend each rationale with “Let’s think in X language” and fix this prompt at inference to control the CoT language. Table 3 reports that on Irish AIME2024 (olympiad‑level), all methods perform similarly, but English‑CoT prompted models clearly outperform on both LC2024 subsets. This suggests that in data‑limited scenarios, directing model to think in English remains the most effective strategy for robust multilingual reasoning.
Figure 5: Performance of English-Pivoted CoT training on Irish AIME and LC2024 with varying number of training samples on r1-distill-Llama-8B (default is NN).
Figure 6: Performance of English-Pivoted CoT training with different $\alpha$ (default: $1.0$) during training ($\beta=1.0$) on r1-distill-Llama-8B. x-axis is evenly spaced for visualization.
@@
r1-distill-Llama-8B Zero-shot Prompting 43.33 79.6 49.74 54.8
r1-distill-Llama-8B Native CoT Training 45.00 82.0 31.90 61.2
@@
DeepHermes-3-Llama-3-8B Zero-shot Prompting 1.67 85.2 3.69 21.6
DeepHermes-3-Llama-3-8B Native CoT Training 3.33 70.8 6.26 44.8
@@
Table 4: Ablation study of fine-tuning on French - a medium resource language, and benchmark across English and French reasoning datasets. Bold scores indicate the best performance per benchmark.
Model Setting AIME2024 (en) MSGM (en) MSGM (zh)
r1-distill-Llama-8B Zero-shot Prompting 43.33 79.6 65.2
r1-distill-Llama-8B Native CoT Training 46.33 92.4 81.6
DeepHermes-3-Llama-3-8B Zero-shot Prompting 1.67 85.2 50.8
DeepHermes-3-Llama-3-8B Native CoT Training 11.67 83.2 61.6
@@
Sample efficiency of English-Pivoted CoT Training. By decoupling reasoning from language learning, our method needs only comprehension fine‑tuning. As Figure 5 shows, with just 1,000 samples ($N/8$) it beats the base model by over 5%, and accuracy climbs further with more data. This highlights its strong performance in low‑resource settings.
Learning dynamic of English-Pivoted CoT Training. Varying the CoT‑loss weight $\alpha$ (with $\beta=1$) in Eq. 5 (Figure 6) reveals that both reducing and over‑emphasizing $\alpha$ degrade performance. This highlights the need to balance the English‑CoT and target‑response objectives. These results underscore the need to balance the reasoning trace and final response objectives in English‑Pivoted CoT Training.
@@
Figure 7: Changes in correction rate from r1-distill-Llama-8B when finetuned with our proposed approach across 3 data resources, low (Irish), medium (French), and high (Chinese).
@@
Generalizability to medium- and high-resource languages. Experiments on Chinese and French - medium to high-resource languages (11234/1-5787; chang-etal-2024-multilinguality), present in Table 4 and Table 5, indicate the generalizability: on French MGSM, r1‑distill‑Llama‑8B scores 54.8%, while our approach boosts it to 83.2%; on Chinese, English pivoting yields only marginal gains over baseline 65.2%, likely because the model naturally reasons in Chinese, so forcing English reasoning can cause interference. Furthermore, the percentage of the new model corrects an incorrect answer by the base model and vice versa are shown in Figure 7. We can see that on MGSM (zh), the correct $\rightarrow$ incorrect rate is the highest, reflecting these conflicts. Nevertheless, our method still achieves an average improvement of 14.30% across the 3 resource regimes (low, medium, and high-resources), demonstrating its broad applicability.
Conclusion
@@
In this work, we propose a novel approach named English-Pivoted CoT Training to effectively transfer the reasoning capabilities of LLMs to low-resource languages. Through training to explicitly aligning reasoning processes across languages by forcing CoT traces to be in the dominant language, our approach achieves significant performance gains (up to 28.33%) over existing techniques, and is generalizable to other resource regimes (medium and high-resource languages). Furthermore, our analysis provides insights into multilingual LLM behavior, particularly the benefit of separating language understanding from reasoning. By contributing the first-ever Irish mathematical reasoning benchmark (LC2024), we also aim to support future research in multilingual reasoning. Future work can be interested in extending to other low-resource languages and tasks, and developing strategies to further enhance cross-lingual reasoning.
Ethics Statement
Our work advances language technologies for low-resource and endangered languages, with a particular focus on Irish. By leveraging English-Pivoted CoT Training, we aim to enhance the accessibility and usability of these languages, thereby promoting linguistic diversity and preserving cultural heritage.
All training and evaluation data are drawn from publicly available sources. Our newly released LC2024 dataset is based on Regulations on Re-use of Public Sector Information 2005, SI 279 of 2005, and is licensed for non-commercial research and educational use only.
Acknowledgments
We would like to acknowledge CloudCIX Limited for the generous collaborative support of computing resources on their NVIDIA HGX/H100 GPU cluster. This publication has emanated from research supported in part by grants from Research Ireland under Grant [12-RC-2289-P2] and [18/CRT/6223] which is co-funded under the European Regional Development Fund. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.